A man rows 12 miles downstream on a river in 2 hrs. He requires another 3 hrs. to row back to his starting position. Find the speed of the current and the rate the man can row in still water.


1
Expert's answer
2022-04-05T01:44:47-0400

"v-rate\\,\\,of\\,\\,man\\\\u-speed\\,\\,of\\,\\,the\\,\\,current\\\\\\left\\{ \\begin{array}{c}\t2\\left( v+u \\right) =12\\\\\t3\\left( v-u \\right) =12\\\\\\end{array} \\right. \\Rightarrow \\left\\{ \\begin{array}{c}\tv+u=6\\\\\tv-u=4\\\\\\end{array} \\right. \\Rightarrow \\left\\{ \\begin{array}{c}\tv=5\\\\\tu=1\\\\\\end{array} \\right."